# Human Image Personalization with High-fidelity Identity Preservation

### v1

@@ 4.3 Benchmark Results
We create a benchmark with the 200 ID clusters in our validation split. We select 4 reference images and 1 target image from each cluster. To evaluate the generated images, we measure the face similarity using a face recognition model (ArcFace-R101 [6]). The most important metric is the face similarity between the generated image and the target image, denoted as $Sim_{Target}$. Additionally, we calculate the maximum face similarity between the generated images and all reference faces, denoted as $Sim_{Ref}$. We also report the CLIP similarity score. To demonstrate that the model can extract ID information and vary following language prompts instead of simply pasting a reference face onto the target image, we calculate $Paste=Sim_{Ref}-Sim_{Target}$ as a metric. A higher value of $Paste$ indicates a greater likelihood of pasting one of the reference images. For the single reference image method, we simply sample the first reference image as input. We also report the FID score [11, 26] as a metric of image quality.
Table 1: Performance comparison of FlashFace and previous methods. We report the face similarity score between the generated image and the target image($Sim_{Target}$), as well as the maximum similarity between the generated face and all reference faces ($Sim_{Ref}$). We also provide the $CLIP$ similarity score of the image. A higher value of $Paste=Sim_{Ref}-Sim_{Target}$ indicates a greater likelihood of pasting one reference face. Our FlashFace achieves the best results in terms of $Sim_{Target}$, demonstrating a high fidelity of ID and better language following ability.
@@
IP-Adaptor(SD1.5) [47] 33.4 57.5 7.7 25.7 121.1
FastComposer(SD1.5) [45] 33.3 65.2 7.5 25.8 104.9
PhotoMaker(SDXL) [21] 45.8 72.1 5.8 40.0 84.2
InstantID(SDXL) [43] 68.6 62.7 16.0 52.6 106.8
@@
As shown in Tab. 1, IP-Adapter [47], FastComposer [45] and PhotoMaker [21] demonstrate a low score for both $Sim_{Ref}$ and $Sim_{Target}$, indicating challenges in preserving facial details when using their embedding-based pipeline. For our concurrent work, InstantID [43], a very high value Paste suggests a strong tendency of reference face pasting instead of following the language prompt. Our method outperforms others in terms of $Sim_{Targe}$ with a reasonable Paste. This suggests that the FlashFace is capable of doing facial variations based on the language prompt while maintaining a high level of identity fidelity.
@@ 4.4 Ablation study
Details preserving ability of ReferenceNet. We assess the unbounded ability of ReferenceNet to maintain face details. We train our model using the cropped target face as the input reference, similar to IP-Adapter [47] and FastComposer [45]. As shown in  Tab. 2, FlashFace can achieve much higher $Sim_{Ref}$ than other methods, indicating the strong ability of ReferenceNet to preserve face details.
@@
IP-Adaptor [47] FastComposer [45] Instant-ID [43] FlashFace
@@
Face reference strength. Fig. 8.(A) reveals a fidelity($Sim_{Target}$) improvement when enhancing the reference strength with $\lambda_{feat}$ and $\lambda_{guidence}$. However, the $Paste$ also increases as shown in Fig. 8.(B). This suggests that there is also an increasing tendency to copy the reference image rather than incorporating language prompts. To achieve a balance between fidelity and language prompting ability, we suggest setting $\lambda_{feat}$ between 0.75 and 1, and $\lambda_{guidence}$ between 2 and 2.5.
Figure 8: The surface plot of $Sim_{Target}$ and $Paste$ with respect to $\lambda_{feat}$ and $\lambda_{guidance}$. Both values, $Paste$ and $Sim_{Target}$, increase as $\lambda_{feat}$ and $\lambda_{guidance}$ increase. To balance fidelity and language prompting ability, we recommend setting $\lambda_{feat}$ between 0.75 and 1, and $\lambda_{guidence}$ between 2 and 2.5.
We also note that $\lambda_{feat}$ exhibits smoother $Sim_{Target}$ changes as it operates in the feature space. We visualize the results for $\lambda_{feat}$ ranging from 0 to 1 under $\lambda_{guidence}=2$ to show the smooth visual similarity improvement (as shown in Fig. 9). Notably, our model consistently produces natural results across the entire $\lambda_{feat}$ range. Another noteworthy observation is that in cases of conflicting conditions, such as when attempting to create a bronze-like face of “Dwayne Johnson”, the optimal result is not achieved at $\lambda_{feat}=1$. At that point, the facial features lose their metallic luster. This highlights an advantage of our framework: the ability to balance conflicts by adjusting the control strength between the text and reference. In our Supplementary Material, we demonstrate the smooth change characteristics of $\lambda_{feat}$ for identity mixing and highlight the role of $\lambda_{guidence}$ in preserving details.
@@
Number of input reference faces. Tab. 3 presents the evaluation results of FlashFace using different numbers of reference images. Increasing the number of reference faces leads to a noticeable improvement in identity enhancement. Specifically, the improvement is significant when changing from 1 reference face to 2, resulting in a 9.3% increase. Another crucial aspect is the reduction of $Paste$. This reduction implies that the model can learn more precise identity information and generate high-quality images based on the prompt, rather than simply relying on reference images.
@@
Tab. 3 presents the evaluation results of FlashFace using different numbers of reference images. Increasing the number of reference faces leads to a noticeable improvement in identity enhancement. Specifically, the improvement is significant when changing from 1 reference face to 2, resulting in a 9.3% increase. Another crucial aspect is the reduction of $Paste$. This reduction implies that the model can learn more precise identity information and generate high-quality images based on the prompt, rather than simply relying on reference images.
@@
The consistent visualization results can be found in Fig. 10, which contains generated images with the prompt “A young woman in a white dress”. Increasing the number of reference images from 1 to 2 indeed significantly enhances the fidelity of identity. When only one reference image is used, the model can capture certain facial features such as the nose and eyes. However, it struggles to accurately depict the overall face shape(The generated face is not round enough). Nevertheless, as the number of reference images increases, the model gradually improves in capturing all these details.
Position of the reference attention layer. As shown in Tab. 4, we conduct ablation experiments by adding reference attention layers to different positions of the U-Net, including the encoder blocks, decoder blocks, or both(Adding reference attention layer to middle blocks by default). We observe that when reference attention layers are only added to the encoder blocks, the model is hard to capture the face identity. This can be attributed to the encoder being further away from the output, making it difficult to align with the convergence direction indicated by the loss. Corresponding to this, when reference attention layers are added solely to the decoder blocks, we achieve the best performance even compared to adding them in both the encoder and decoder.

### v2

@@ 4.3 Benchmark Results
We create a benchmark with the 200 ID clusters in our validation split. We select 4 reference images and 1 target image from each cluster. To evaluate the generated images, we measure the face similarity using a face recognition model (ArcFace-R101 [Deng_2022]). The most important metric is the face similarity between the generated image and the target image, denoted as $Sim_{Target}$. Additionally, we calculate the maximum face similarity between the generated images and all reference faces, denoted as $Sim_{Ref}$. We also report the CLIP similarity score. To demonstrate that the model can extract ID information and vary following language prompts instead of simply pasting a reference face onto the target image, we calculate $Paste=Sim_{Ref}-Sim_{Target}$ as a metric. A higher value of $Paste$ indicates a greater likelihood of pasting one of the reference images. For the single reference image method, we simply sample the first reference image as input. We also report the FID score [heusel2018gans, parmar2022aliased] as a metric of image quality.
Table 1: Performance comparison of FlashFace and previous methods. We report the face similarity score between the generated image and the target image($Sim_{Target}$), as well as the maximum similarity between the generated face and all reference faces ($Sim_{Ref}$). We also provide the CLIPCLIP similarity score of the image. A higher value of $Paste=Sim_{Ref}-Sim_{Target}$ indicates a greater likelihood of pasting one reference face. Our FlashFace achieves the best results in terms of $Sim_{Target}$, demonstrating a high fidelity of ID and better language following ability.
@@
IP-Adaptor(SD1.5) [ye2023ip-adapter] 33.4 57.5 7.7 25.7 121.1
FastComposer(SD1.5) [xiao2023fastcomposer] 33.3 65.2 7.5 25.8 104.9
PhotoMaker(SDXL) [li2023photomaker] 45.8 72.1 5.8 40.0 84.2
InstantID(SDXL) [wang2024instantid] 68.6 62.7 16.0 52.6 106.8
@@
As shown in Tab.˜1, IP-Adapter [ye2023ip-adapter], FastComposer [xiao2023fastcomposer] and PhotoMaker [li2023photomaker] demonstrate a low score for both $Sim_{Ref}$ and $Sim_{Target}$, indicating challenges in preserving facial details when using their embedding-based pipeline. For our concurrent work, InstantID [wang2024instantid], a very high value Paste suggests a strong tendency of reference face pasting instead of following the language prompt. Our method outperforms others in terms of $Sim_{Targe}$ with a reasonable Paste. This suggests that the FlashFace is capable of doing facial variations based on the language prompt while maintaining a high level of identity fidelity.
@@ 4.4 Ablation study
Details preserving ability of ReferenceNet. We assess the unbounded ability of ReferenceNet to maintain face details. We train our model using the cropped target face as the input reference, similar to IP-Adapter [ye2023ip-adapter] and FastComposer [xiao2023fastcomposer]. As shown in  Tab.˜2, FlashFace can achieve much higher $Sim_{Ref}$ than other methods, indicating the strong ability of ReferenceNet to preserve face details.
@@
IP-Adaptor [ye2023ip-adapter] FastComposer [xiao2023fastcomposer] Instant-ID [wang2024instantid] FlashFace
@@
Face reference strength. Fig.˜8.(A) reveals a fidelity($Sim_{Target}$) improvement when enhancing the reference strength with $\lambda_{feat}$ and $\lambda_{guidence}$. However, the $Paste$ also increases as shown in Fig.˜8.(B). This suggests that there is also an increasing tendency to copy the reference image rather than incorporating language prompts. To achieve a balance between fidelity and language prompting ability, we suggest setting $\lambda_{feat}$ between 0.75 and 1, and $\lambda_{guidence}$ between 2 and 2.5.
Figure 8: The surface plot of $Sim_{Target}$ and PastePaste with respect to $\lambda_{feat}$ and $\lambda_{guidance}$. Both values, PastePaste and $Sim_{Target}$, increase as $\lambda_{feat}$ and $\lambda_{guidance}$ increase. To balance fidelity and language prompting ability, we recommend setting $\lambda_{feat}$ between 0.75 and 1, and $\lambda_{guidence}$ between 2 and 2.5.
We also note that $\lambda_{feat}$ exhibits smoother $Sim_{Target}$ changes as it operates in the feature space. We visualize the results for $\lambda_{feat}$ ranging from 0 to 1 under $\lambda_{guidence}=2$ to show the smooth visual similarity improvement (as shown in Fig.˜9). Notably, our model consistently produces natural results across the entire $\lambda_{feat}$ range. Another noteworthy observation is that in cases of conflicting conditions, such as when attempting to create a bronze-like face of “Dwayne Johnson”, the optimal result is not achieved at $\lambda_{feat}=1$. At that point, the facial features lose their metallic luster. This highlights an advantage of our framework: the ability to balance conflicts by adjusting the control strength between the text and reference. In our Supplementary Material, we demonstrate the smooth change characteristics of $\lambda_{feat}$ for identity mixing and highlight the role of $\lambda_{guidence}$ in preserving details.
@@
Number of input reference faces. Tab.˜3 presents the evaluation results of FlashFace using different numbers of reference images. Increasing the number of reference faces leads to a noticeable improvement in identity enhancement. Specifically, the improvement is significant when changing from 1 reference face to 2, resulting in a 9.3% increase. Another crucial aspect is the reduction of $Paste$. This reduction implies that the model can learn more precise identity information and generate high-quality images based on the prompt, rather than simply relying on reference images.
@@
The consistent visualization results can be found in Fig.˜10, which contains generated images with the prompt “A young woman in a white dress”. Increasing the number of reference images from 1 to 2 indeed significantly enhances the fidelity of identity. When only one reference image is used, the model can capture certain facial features such as the nose and eyes. However, it struggles to accurately depict the overall face shape(The generated face is not round enough). Nevertheless, as the number of reference images increases, the model gradually improves in capturing all these details.
Position of the reference attention layer. As shown in Tab.˜4, we conduct ablation experiments by adding reference attention layers to different positions of the U-Net, including the encoder blocks, decoder blocks, or both(Adding reference attention layer to middle blocks by default). We observe that when reference attention layers are only added to the encoder blocks, the model is hard to capture the face identity. This can be attributed to the encoder being further away from the output, making it difficult to align with the convergence direction indicated by the loss. Corresponding to this, when reference attention layers are added solely to the decoder blocks, we achieve the best performance even compared to adding them in both the encoder and decoder.